OpenStudy (anonymous):

given the circle x^2 + y^2 = 16, if you translate the circle 5 units to the left and 4 units down, what is the new equation?

OpenStudy (anonymous):

go to https//webmath.com//

OpenStudy (anonymous):

(x-a)^2 + (y-b)^2 = r^2 is the equation of a circle, where r is the radius of the circle, and the center of the circle is (a, b) Looking at your original equation: x^2 + y^2 = 16, the radius is sqrt(16) = 4 and the center is (0,0) What is the new center? Use that to write your new equation.